Find the area contained between three circles of radius 20 centimeters, each of which is externally tangent to the other two.
Answer by sachi(548)

the circles are externally tangent to the other two.
the lines joining the centers of the circles forms an equilateral triangle with sides = 2* radius=40 cms & the area of the triangle= (sqrt3/4)*40^2=692.8 cm sqr
the two radii of the circle form a sector with angl 60 degree
the area of the 3 sectors=3*(60/360)(pi radius^2) =3/6*22/7*20^2=628.57 cm sqr
the area contained between three circles of radius
=the area of the triangle-the area of the 3 sectors
=692.8-628.57=64.23cm sqr
ans